OpenStudy (anonymous):

(tan^2 θ − 16)(2 cos θ + 1) = 0

OpenStudy (tkhunny):

Have you considered factoring the difference of squares on the tangent piece? Then you're off!

OpenStudy (anonymous):

how do you do that?

OpenStudy (tkhunny):

Factor this expression: x^2 - y^2

OpenStudy (anonymous):

(x + y)(x - y)

OpenStudy (tkhunny):

Perfect. Do the exact same thing with this expression: (tan^2 θ − 16)

OpenStudy (anonymous):

\[(\tan \theta +4)(\tan \theta-4)\]

OpenStudy (tkhunny):

Super. \((tan(\theta) − 4)(tan(\theta) + 4)(2 cos θ + 1) = 0\) Now, what do you do with a fully factored expression set to zero to find all the possible solutions?

OpenStudy (anonymous):

so \[\tan \theta+4=0? \tan \theta-4=0? 2\cos+1=0?\]

OpenStudy (tkhunny):

You have it!! Track down those solutions. There are lots.

OpenStudy (anonymous):

i hsvr twenty of these if youre willing to work these out with me

OpenStudy (tkhunny):

Did you get this one? I haven't seen the results, yet.

OpenStudy (anonymous):

1.325 -1.325 pi/3

OpenStudy (anonymous):

are those all of the solutions?

OpenStudy (tkhunny):

Yes, plus many more, since tangent and cosine are periodic. 2cos(x) + 1 = 0 cos(x) = -1/2 x = 2pi/3 + 2kpi AND 4pi/3 + 2kpi for any integer, k. That's quite a few solutions - infinitely many! Do the same thing with the tangents.

OpenStudy (anonymous):

Let k be any integer. Round terms to three decimal places where appropriate. If there is no solution, enter NO SOLUTION.) those are the contraints added to the problem

OpenStudy (tkhunny):

There you go. I did it for the cosine piece. You have an initial answer for two tangent pieces. You will need to find the rest.

OpenStudy (tkhunny):

Tangent is positive in 1st and 3rd quadrants.

OpenStudy (anonymous):

so its x= pi/3+kpi and 5pi/3+kpi??

OpenStudy (anonymous):

what my problem is that im not understanding the fundamental concepts. I need something to explain how each of these work

OpenStudy (tkhunny):

No, cos(x) = -1/2 x = 2pi/3 or 4pi/3 You're in the wrong quadrants.

OpenStudy (anonymous):

ohh I was in quadrents 1 and 4 so are those the only two for cos? so that makes the tan answers: only 1.325 and -1.325

OpenStudy (tkhunny):

No, there are lots more to go. You have x = 1.325. What is the period of the tangent function?

OpenStudy (anonymous):

im guessing its 2pi it doesnt say

OpenStudy (tkhunny):

Good guess, but it's just pi. You're almost done!

OpenStudy (anonymous):

so 1.325, and 2pi/3 4pi/3

OpenStudy (tkhunny):

No, where did the k go and where did all the other infinitely many solutions go? \(\theta = 2\pi/3 + 2k\pi\) \(\theta = 4\pi/3 + 2k\pi\) \(\theta = 1.325 + k\pi\) \(\theta = -1.325 + k\pi\) That's a whole lot of answers!